# Quantum Theory and Unusual Dielectric Functions of Graphene

### v1

@@ I Introduction
The response functions of graphene were investigated in both the spatially local (${\bf q}=0,\varepsilon^{\rm L}=\varepsilon^{\rm T}=\varepsilon$) and nonlocal cases using a number of more or less phenomenological approaches such as the hydrodynamic model 5a ; 5b ; 5c , Boltzmann transport theory and the Drude model 11 ; 11a ; 11b ; 13 ; 14 ; 18 ; 24 ; 25 ; 31 , current-current correlation functions and the random phase approximation 9 ; 12 ; 12a ; 18a ; 21 ; 23a ; 26 ; 26a ; 26b ; 29a , density-functional theory 23 ; 23aa , Kubo response theory 6 ; 7 ; 7a ; 7ab ; 8 ; 15 ; 16 ; 17 ; 27 ; 27a ; 28 ; 29 , modeling graphene optics in terms of Lorentz-type oscillators 31a and by using the Fresnel reflection coefficients 32 , etc. (see also the reviews 34 ; 35 ; 36 ). The obtained results are of different levels of accuracy and areas of application. In the framework of the Dirac model, however, the response functions of graphene can be found exactly starting from the first principles of thermal quantum field theory using the formalism of the polarization tensor in (2+1)-dimensions.
@@ Appendix A The list of auxiliary functions used in the article
$\alpha$ $\frac{e^{2}}{\hbar c}$ — fine structure constant

### v2

@@ I Introduction
The response functions of graphene were investigated in both the spatially local ($\mbox{$q$}=0,~\varepsilon^{\rm L}=\varepsilon^{\rm T}=\varepsilon$) and nonlocal cases using a number of more or less phenomenological approaches such as the hydrodynamic model 5a ; 5b ; 5c , Boltzmann transport theory and the Drude model 11 ; 11a ; 11b ; 13 ; 14 ; 18 ; 24 ; 25 ; 31 , current-current correlation functions and the random phase approximation 9 ; 12 ; 12a ; 18a ; 21 ; 23a ; 26 ; 26a ; 26b ; 29a , density-functional theory 23 ; 23aa , Kubo response theory 6 ; 7 ; 7a ; 7ab ; 8 ; 15 ; 16 ; 17 ; 27 ; 27a ; 28 ; 29 , modeling graphene optics in terms of Lorentz-type oscillators 31a and by using the Fresnel reflection coefficients 32 , etc. (see also the reviews 34 ; 35 ; 36 ). The obtained results are of different levels of accuracy and areas of application. In the framework of the Dirac model, however, the response functions of graphene can be found exactly starting from the first principles of thermal quantum field theory using the formalism of the polarization tensor in (2+1)-dimensions.
@@ VI Conclusions
This work was supported by the State Assignment for Basic Research (project FSEG-2026-0018).
@@ Appendix A The list of auxiliary functions used in the article
$\alpha$ $\frac{e^{2}}{\hbar c}$ is the fine structure constant